OpenStudy (anonymous):

URGENT. If 2^(x+4) = 3^x. Prove by contradiction that x is an irrational number.

OpenStudy (accessdenied):

Well, I think if we are trying to prove something with a contradiction, we should start by letting x be what it shouldn't be, and then searching for a logical failure...

OpenStudy (anonymous):

I started with x being a/b which is rational form, but i dont know how to prove that its not.

OpenStudy (accessdenied):

So, if we plug that in, we get... \(\displaystyle 2^{a/b + 4} = 3^{a/b} \)

OpenStudy (anonymous):

yeah, so the only operations that i found that we could do would be base rooting, but that doesnt seem to help

OpenStudy (accessdenied):

We could take the logarithm of both sides here. \( \displaystyle \log \left( 2^{a/b + 4} \right) = \log \left(3^{a/b} \right) \) \(\displaystyle (\frac{a}{b} + 4) \log 2 = \frac{a}{b} \log 3 \)

OpenStudy (accessdenied):

Well, I think if we get a/b on one side of the equality, there should be some expression of logarithms on the other side.

OpenStudy (anonymous):

thanks for your help btw

OpenStudy (accessdenied):

We could do that, although I think it'd be easier even like this: \( \displaystyle \frac{a}{b} \log 2 + 4 \log 2 = \frac{a}{b} \log 3 \) and we subtract \( \frac{a}{b} \log 3 \) from both sides and \(4\log 2 \) from both sides. We can then factor out an a/b and divide... \( \displaystyle \frac{a}{b} \left( \log 2 - \log 3 \right) = - 4 \log 2 \) \( \displaystyle \frac{a}{b} = \frac{ -4 \log 2 }{ \log 2 - \log 3} \)

OpenStudy (accessdenied):

I think from there, its pretty clear that our right-hand side is not rational... -4log 2 and log2 - log3 are not integers, so their quotient would not be rational.

OpenStudy (accessdenied):

You're welcome! :)

OpenStudy (anonymous):

hey can you do one more thing can you just delete your messages with the answers in them? thanks man

OpenStudy (anonymous):

this problem isnt for anything important, but sometimes online help is looked down upon for homework